OpenStudy (anonymous):

Solve the equation. 4(y + 3) = 28

myininaya (myininaya):

distribute the 4 4y+12=29 subtract 12 4y+12-12=29-12 4y=17 divide by 4 y=17/4

OpenStudy (anonymous):

4(y + 3) = 28 (y + 3) = 28/4 y= 7-3 =4

myininaya (myininaya):

lol thats if you have 4(y+3)=29 i gave you an example to this problem 4(y+3)=28 same process k?

myininaya (myininaya):

4y+12=28 4y=28-12 4y=16 y=4 you could have also done as mark done above divide by 4 on both sides first

OpenStudy (anonymous):

oh thanks